OpenStudy (anonymous):

What is the mean of the following numbers: 2, 4, 3, 4, 4, 1, 3

OpenStudy (anonymous):

Sum them a divide by 7

OpenStudy (anonymous):

\[ \frac 1 7 ( 2+4+3+4+ 4+1+ 3)= \frac {21}7=3 \]